OpenStudy (anonymous):

Cammy wrote the expression 8a+6(b-2a) . Fred wrote an expression equivalent to cammy. Which expression could Fred have written. 6b-4a 20a+6b 20a-6b 4a+6b

OpenStudy (instagrammodel):

Use the distributive property: 8a + 6(b - 2a) = 8a + 6(b) - 6(2a) = 8a + 6b - 12a = <--- Finish this to get the answer!